order; Fortaz 250mg IM. The vial contains 500 mg powder with directions to add 1.5 ml of sterile water for a total volume of 1,8 mL. How many mL will you administer?
Answer #1

ANSWER.

* Calculation of amount of Fortaz needed for administration.

             Formula = Desired dose / Dose in hand * Quantity.

                   Desired dose (D) = 250 mg.

                   Dose in hand (H) = 500 mg.

                            Quantity (Q) = 1.8 ml.

           As per formula

                Amount (ml) of drug required for administration = D/H *Q

                                                                                            = 250/500 * 1.8

                                                                                            = 0.5 *1.8

                                                                                            = 0.9 ml.